(a.) Light of wavelength 5.5 × 10-7m falls on a single silt of width 0.15mm. A screen is placed 1.2m beyond the slit.

             (i.) Name the phenomenon that occurs.

              (ii) Sketch a graph showing the light pattern observed on the screen.

             (iii.) Calculate the width of the central fringe.



1
Expert's answer
2021-02-22T11:08:18-0500

(I). The phenomenon is diffraction.

(II).


(III).

"Width\:of\:central\:fringe\left(Y_m\right)"

"=\frac{Order\:of\:maximum\left(m\right)\times Wavelength\left(\lambda \right)\times distance\:from\:slit\left(L\right)}{Width\:of\:the\:slit\left(D\right)}"

"=\frac{m\lambda L}{D}"

"=\frac{\left(1x5.5x10^{-7}x1.2\right)}{\left(0.15x10^{-3}\right)}"

"=4.4x10^{-3}m"